# Maximal work extraction unitarily from an unknown quantum state: Ergotropy estimation via feedback experiments

### v1

@@ IV.2 Two-qubit ergotropy
To demonstrate local and global ergotropy estimation, we prepare five initial states with varying degrees of entanglement between the system qubits $\mathrm{H}$ and $\mathrm{F}$. The full NMR pulse-sequences are described in the Appendix. After preparing $|{000}\rangle\langle{000}|$ we transform it to $|{0}\rangle\langle{0}|\otimes\rho^{\mathrm{HF}}_{0}=|{0}\rangle\langle{0}|\otimes U_{G}|{00}\rangle\langle{00}|U_{G}^{\dagger}$ using a global unitary $U_{G}$ consisting of a Hadamard operator on $\mathrm{H}$ and a controlled $e^{-i\nu I_{x}}$ gate on $\mathrm{F}$. Note that the parameter $\nu$ controls the degree of entanglement. That completes the initialization, and now we start work extraction by FQErgo. During the first 10 iterations, we extract work by using only local drives $I_{x(y)}^{\mathrm{H}}$, $I_{x(y)}^{\mathrm{F}}$ (similar as in Fig. 3 (a)) and realize the first energy minimization corresponding to the local passive state as shown in Fig. 5 (c). From 11th to 20th iterations, we extract further energy via the global drive $I_{x}^{\mathrm{H}}I_{y}^{\mathrm{F}}+I_{y}^{\mathrm{H}}I_{x}^{\mathrm{F}}$ along with the $Z_{\delta}$ gate and local drives as shown in Fig. 3 (b). The global drive allows complete work extraction, eventually taking the system qubits to the second energy minimization corresponding to the global passive state $\rho^{p}_{\mathrm{HF}}$. Fig. 5 (b) shows the normalized energies for both local and global extraction with initial states having varying entanglement entropy $S=-\mathrm{Tr}[\rho^{\mathrm{H}}_{0}\log\rho^{\mathrm{H}}_{0}]$, where $\rho^{\mathrm{H}}_{0}=\mathrm{Tr}_{\mathrm{F}}(\rho_{0}^{\mathrm{HF}})$. Fig. 5 (c-d) plot the experimentally estimated local and global ergotropy values vs simulated values and Fig. 5 (e) plots the experimentally obtained ergotropy gaps vs versus the simulated values. The good agreement between the estimated values and the simulated values confirms the successful demonstration of the FQErgo algorithm.
@@ V Summary and discussions
We envisage several future directions. For instance, previously using ergotropy as an entanglement witness required prior knowledge of the class of states (eg. [41]). FQErgo can overcome such limitations and pave the way to certify the entanglement for a completely unknown state. Since one probe-qubit suffices, irrespective of the system size, extending FQErgo to larger systems should be feasible without exponential complexity. Although the ensemble nature of NMR is advantageous here, the overall algorithm is general enough to adapt to other quantum architectures.
@@ Acknowledgements
Figure 6: NMR pulse sequence for the 1-qubit FQErgo experiment. From thermal equilibrium state, we first prepare the $|{00}\rangle$ pseudopure state (PPS) which is rotated by $\theta_{\phi}$ to realize the initial state $\rho(0)$. After FQErgo iterations $U_{d}(\beta)$ on the system qubit 31P (only one iteration is shown) and Hadamard gate (H) on the probe, we extract (i) one of the drive amplitudes $\beta_{x(y)}$ or (ii) energy $E$. Here the measurement pulse-sequences implement one of $e^{-i\beta^{x/y/z}H_{x/y/z}\tau}$ acting on system 31P qubit and controlled by 19F qubit. Finally, $\sigma_{x}$ measurements of 19F qubit yields drive amplitudes $\beta_{x(y)}$ or energy $E$.
Figure 7: The NMR pulse sequences for 2-qubit (1H,19F), FQEergo with one (13C) probe qubit: (a) PPS preparation, initializing the system as $\rho_{23}(0)$ using entangling unitary $U_{G}$ after a pseudo-Hadamard on 1H. Here delay $d\tau$ is varied to change entanglement entropy. (b) Local and global drives, where $UXY=e^{-i\beta^{xy}H_{dxy}\tau}$ and the delay $d\beta$ controls the strength of the global drive. (c) Extracting global drive strength $\beta^{xy}$ and system energy.

### v2

@@ IV.2 Two-qubit ergotropy
To demonstrate local and global ergotropy estimation, we prepare five initial states with varying degrees of entanglement between the system qubits $\mathrm{H}$ and $\mathrm{F}$. The full NMR pulse-sequences are described in the Appendix. After preparing $|{000}\rangle\langle{000}|$ we transform it to $|{0}\rangle\langle{0}|\otimes\rho^{\mathrm{HF}}_{0}=|{0}\rangle\langle{0}|\otimes U_{G}|{00}\rangle\langle{00}|U_{G}^{\dagger}$ using a global unitary $U_{G}$ consisting of a Hadamard operator on $\mathrm{H}$ and a controlled $e^{-i\nu I_{x}}$ gate on $\mathrm{F}$, where $\nu$ controls the degree of entanglement. That completes the initialization, and now we start work extraction by FQErgo. During the first 10 iterations, we extract work by using only local drives $I_{x(y)}^{\mathrm{H}}$, $I_{x(y)}^{\mathrm{F}}$ (similar as in Fig. 3 (a)) and realize the first energy minimization corresponding to the local passive state as shown in Fig. 5 (c). From 11th to 20th iterations, we extract further energy via the global drive $I_{x}^{\mathrm{H}}I_{y}^{\mathrm{F}}+I_{y}^{\mathrm{H}}I_{x}^{\mathrm{F}}$ along with the $Z_{\delta}$ gate and local drives as shown in Fig. 3 (b). The global drive allows complete work extraction, eventually taking the system qubits to the second energy minimization corresponding to the global passive state $\rho^{p}_{\mathrm{HF}}$. Fig. 5 (b) shows the normalized energies for both local and global extraction with initial states having varying entanglement entropy $S=-\mathrm{Tr}[\rho^{\mathrm{H}}_{0}\log\rho^{\mathrm{H}}_{0}]$, where $\rho^{\mathrm{H}}_{0}=\mathrm{Tr}_{\mathrm{F}}(\rho_{0}^{\mathrm{HF}})$. Fig. 5 (c-d) plot the experimentally estimated local and global ergotropy values vs simulated values and Fig. 5 (e) plots the experimentally obtained ergotropy gaps vs versus the simulated values. The good agreement between the estimated values and the simulated values confirms the successful demonstration of the FQErgo algorithm.
@@ V Summary and discussions
We envisage several future directions. For instance, while ergotropy based thermodynamic quantities can form convenient entanglement witnesses, they required prior knowledge of the class of states (eg. [43]). FQErgo can overcome such limitations and pave the way to certify the entanglement for a completely unknown state. Since one probe-qubit suffices, irrespective of the system size, extending FQErgo to larger systems should be feasible without exponential complexity. The procedure for extracting work from an unknown state is crucial from the perspective of quantum batteries. Although a fully charged quantum battery may start from a known state, subsequently during storage or partial usage it may end up in an unknown state. Although the ensemble nature of NMR is advantageous here, the overall algorithm is general enough to adapt to other quantum architectures.
@@ Acknowledgements
Figure 6: NMR pulse sequence for the 1-qubit FQErgo experiment. From thermal equilibrium state, we first prepare the $|{00}\rangle$ pseudopure state (PPS) which is rotated by $U_{\theta,\phi}=e^{-i\theta(I_{x}\cos\phi+I_{y}\sin\phi)}$ to realize the initial state $\rho_{0}$. After FQErgo iterations $u_{k}=e^{-i\beta^{x}_{k}I_{x}\tau}e^{-i\beta^{y}_{k}I_{y}\tau}$ on the system qubit 31P (only one iteration is shown) and Hadamard gate (H) on the probe, we extract (i) one of the drive amplitudes $\beta^{x(y)}$ or (ii) energy $E$. Here the measurement pulse-sequences implement one of $e^{-i\beta^{x(y)}I_{y(x)}\tau}$ and $e^{-iI_{z}\tau}$ acting on system 31P qubit and controlled by 19F qubit. Finally, $\sigma_{x}$ measurements of 19F qubit yields drive amplitudes $\beta^{x(y)}$ or energy $E$.
Figure 7: The NMR pulse sequences for 2-qubit (1H,19F), FQEergo with one (13C) probe qubit: (a) PPS preparation, initializing the system as $\rho^{\mathrm{HF}}_{0}$ using entangling unitary $U_{G}$ after a pseudo-Hadamard on 1H. Here delay $d\tau$ is varied to change entanglement entropy. (b) Local and global drives, where $\delta_{y}=e^{-i\delta I^{F}_{y}}$, $U_{xy}=e^{-i\beta^{xy}H_{xy}\tau_{xy}}$ and the delay $d\beta$ controls the strength of the global drive. (c) Extracting global drive strength $\beta^{xy}$ and system energy.